A rectangle carpet has perimeter of 234 in. Length is 71 in more than the width what are the dementions of the carpet
Answer by Cromlix(4381)

Perimeter = 2 x length + 2 x width
Width = x
Length = x + 71
234 = 2*(x + 71) + 2*(x)
234 = 2x + 142 + 2x
234 = 4x + 142
4x = 234 - 142
4x = 92
x = 23
Width = 23 inches
Length = 94 inches
Hope this helps.
:-)